# DeepSelective: Feature Gating and Representation Matching for Interpretable Clinical Predictions

### v1

@@ 2.1 Feature Selection for EHRs
Feature selection is crucial for building predictive models from EHRs, traditionally relying on labor-intensive expert-defined features[9]. Deep learning has shifted this paradigm by enabling automatic feature construction, but often lacks interpretability [10]. However, deep learning models involve complex linear and nonlinear transformations, which typically results in poor interpretability [11]. On the contrary, feature selection step ensures the interpretability of a machine learning model by explicitly recommending biologically meaningful biomarkers for training the prediction models [12]. Recent approaches, such as Gronsbell et al.’s automated feature selection method [9] and Liu et al.’s EnRank algorithm [13], aim to balance accuracy with interpretability by integrating biologically meaningful biomarkers.
@@ 2.3 Interpretable Deep Learning Approaches
To enhance the interpretability of deep learning models, some studies integrate traditional interpretable modules [21]. For instance, the wide & deep learning framework combines linear models with deep neural networks to improve both prediction accuracy and interpretability [22]. Other methods, such as NeuralFS and Cancelout, focus on feature selection within neural networks, though they don’t fully leverage deep learning’s representation learning capabilities [23].
@@ 6 Conclusion
In this study, we introduce DeepSelective, a novel framework designed to enhance predictive modeling in healthcare through robust feature selection and compression techniques. The framework integrates the Dynamic Gating Feature Selection (DGFS) module, the Attentive Transformer Autoencoder (ATA), and a SparsityController to achieve a delicate balance between interpretability and performance. Our approach leverages the powerful representational learning capabilities of deep learning while ensuring that only the most relevant features are retained, thereby improving both the accuracy and interpretability of the model.
The experimental results on mortality and decompensation prediction tasks demonstrate the superiority of DeepSelective compared to various baseline models. DeepSelective consistently achieved higher AUROC, AUPRC, and min(Se,P+) scores, validating the effectiveness of our dual-module architecture. The further underscored the importance of each component, with significant performance drops observed when any module was removed. These findings affirm our hypothesis that combining DGFS and ATA modules with a dynamic sparsity control mechanism significantly enhances the model’s ability to handle high-dimensional EHRs data. Despite these promising results, our study has some limitations. One notable limitation is the reliance on predefined temperature ranges for the SparsityController, which requires careful tuning to balance feature sparsity and model performance. Additionally, while DeepSelective demonstrates strong performance across the datasets tested, further validation on more diverse datasets and clinical settings is necessary to generalize its applicability. Future work could explore adaptive mechanisms for temperature tuning and extend the framework to incorporate additional types of clinical data.
In conclusion, DeepSelective represents an advancement in predictive modeling for healthcare, offering a robust and interpretable approach to feature selection and compression. Our comprehensive experimental evaluation highlights its potential to improve clinical decision-making processes, provide a useful tool for more effective and transparent AI-driven healthcare solutions.

### v2

# DeepSelective: Interpretable Prognosis Prediction via Feature Selection and Compression in EHR Data

@@ Abstract
[…]
[e]organization=School of Science,addressline=The Hong Kong University of Science and Technology, city=Hong Kong, postcode=999077, country=China
@@ 2.1 Feature Selection for EHRs
Feature selection is crucial for building predictive models from EHRs, traditionally relying on labor-intensive expert-defined features [20]. Deep learning has shifted this paradigm by enabling automatic feature construction 11, but often lacks interpretability [24]. On the contrary, feature selection step ensures the interpretability of a machine learning model by explicitly recommending biologically meaningful biomarkers for training the prediction models [21]. Recent approaches, such as Gronsbell et al.’s automated feature selection method [20] and Liu et al.’s EnRank algorithm [27], aim to balance accuracy with interpretability by integrating biologically meaningful biomarkers.
@@ 2.3 Interpretable Deep Learning Approaches
To enhance the interpretability of deep learning models, some studies integrate traditional interpretable modules [4]. For instance, the wide & deep learning framework combines linear models with deep neural networks to improve both prediction accuracy and interpretability [40]. Other methods, such as NeuralFS and Cancelout, focus on feature selection within neural networks, though they don’t fully leverage deep learning’s representation learning capabilities [41]. More recently, ISP-IRLNet proposes a joint optimization framework that combines an interpretable sampling module with an implicit regularization network, showing that interpretability can also be embedded in data acquisition and reconstruction processes, further broadening the scope of interpretable deep learning[12].
@@ 6 Conclusion
In this work, we presented DeepSelective, a novel end-to-end deep learning framework tailored for clinical prognosis prediction using high-dimensional EHR data. The framework introduces a dual compression design: a Dynamic Gating Feature Selection (DGFS) module for sparse, interpretable feature selection, and an Attentive Transformer Autoencoder (ATA) for perceptual compression via deep representation learning. Combined with a Representation Matching Layer (RML) and a SparsityController, DeepSelective effectively balances model performance and interpretability, two often conflicting goals in clinical AI applications.
Our extensive experiments on two benchmark clinical prediction tasks—mortality prediction and real-time decompensation detection—demonstrate that DeepSelective consistently outperforms state-of-the-art baselines across multiple metrics (AUROC, AUPRC, min(Se,P+)). Furthermore, the ablation studies confirm the necessity of each component within the framework, showing clear performance degradation when any module is removed. We also validated the clinical meaningfulness of selected features through mutual information analysis, T-tests, and PCA visualizations, highlighting the model’s transparency and relevance to real-world clinical decision-making.
However, our framework has limitations. First, the temperature tuning in the Gumbel-Softmax-based SparsityController is currently guided by heuristic initialization and PID-based adjustments, which might not generalize optimally across all data distributions. Second, although our evaluation on MIMIC-III datasets is thorough, external validation on broader EHR systems—across hospitals, disease domains, and demographic groups—remains to be conducted. In addition, model performance could potentially degrade in the presence of severe missingness or noisy features not addressed by preprocessing alone.
We believe DeepSelective contributes a principled and practical approach to interpretable prognosis modeling. It offers modular components (e.g., plug-in feature selectors, matching layers) that other researchers can adapt for various structured data prediction problems beyond healthcare, such as finance or manufacturing. Moreover, clinicians and data scientists can benefit from its transparent prediction process, where key features and representations are explicitly surfaced. In future work, we plan to develop self-adaptive sparsity control mechanisms using reinforcement learning or meta-learning to replace manual tuning, incorporate multi-modal clinical data such as medical imaging and unstructured notes to improve representation learning, and explore the use of counterfactual explanations to further enhance model interpretability.